OpenStudy (anonymous):

Complete the square to solve the equation below. x2 + 10x - 13 = 17

OpenStudy (anonymous):

ok i will go through it with you: add 13 to both sides. tell me what the resulting equation is

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

what do you get when you add 13 to both sides?

OpenStudy (anonymous):

ok so when you add 13 to both sides you get x2+10x = 30

OpenStudy (anonymous):

then you add half of the b term squared to both sides

OpenStudy (anonymous):

the b term is 10 so half of it is 5 and that squared is 25

OpenStudy (anonymous):

if you add 25 to both sides you get x2+10x+25=55

OpenStudy (anonymous):

the reason we added 25 is now we can write the left hand sdes as (x+half the b term)^2

OpenStudy (anonymous):

half the b term is 5 so we can write th left hand side as (x+5)^2 = 55

OpenStudy (anonymous):

that means x+5 = sqrt(55) or x+5 = -sqrt(55)

OpenStudy (anonymous):

so x equals sqrt(55)-5 or -sqrt(55)-5

OpenStudy (anonymous):

thats it